Solve this confidence interval problem-- Since we are currently conducting a study on millennials and their TV time during the week, I decided to estimate the population mean number of hours that millennials watch more than 5 hours a day and if so, how many hours did they spend watching TV. According to the survey, I pulled 10 individuals randomly and these were the results—

6, 10, 3, 0, 4, 8, 2, 1, 1, 8

Calculate a confidence level of 95%.
Answer #1

Solution:

Given that,

x x2
6 36
10 100
3 9
0 0
4 16
8 64
2 4
1 1
1 1
8 64
\sum x = 43 \sum x2 = 295

The sample mean is \bar x

Mean \bar x  =  \sum x / n

= 43 / 10

= 4.3

Mean \bar x  = 4.3

The sample standard deviation is S

S = \sqrt{} ( \sumx2 ) - (( \sumx )2 / n ) / 1 -n )

= \sqrt{} ( 295 - ( 43 )2 / 10 ) 9

=  \sqrt{} ( 295 - 184 / 9 )

   =  \sqrt{}110.1 / 9

   =  \sqrt{}12.2333

= 3.4976

The sample standard deviation = 3.5

\bar x = 4.3

s = 3.5

n = 10

Degrees of freedom = df = n - 1 = 10- 1 = 9

At 95% confidence level the z is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

t\alpha /2,df = t0.025,9 =2.262

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.262 * (3.5 / \sqrt10)

= 2.504

The 95% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

4.3 - 2.504 < \mu < 4.3 + 2.504

1.796 < \mu < 6.804